### v4

# The Face of Populism: Examining Differences in Facial Emotional Expressions of Political Leaders Using Machine Learning

Abstract
Populist rhetoric employed on online media is characterized as deeply impassioned and often imbued with strong emotions. The aim of this paper is to empirically investigate the differences in affective nonverbal communication of political leaders. We use a deep-learning approach to process a sample of 220 YouTube videos of political leaders from 15 different countries, analyze their facial expressions of emotion and then examine differences in average emotion scores representing the relative presence of 6 emotional states (anger, disgust, fear, happiness, sadness, and surprise) and a neutral expression for each frame of the YouTube video. Based on a sample of manually coded images, we find that this deep-learning approach has 53-60% agreement with human labels. We observe statistically significant differences in the average score of negative emotions between groups of leaders with varying degrees of populist rhetoric.
1 Introduction
In the growing body of literature on populism, the category of emotion has most frequently been utilized in discrete ways to describe its emergence and tenacity (Demertzis 2006)—by reference to the disillusioned silent majority, heightened emotions surrounding political leaders, and emotionally charged divisions between the people and the elite (Moffitt & Tormey 2014, Mudde 2017, Norris & Inglehart 2019). Compared to pluralist parties, who are commonly portrayed as more affectively neutral, populist rhetoric is most often characterized as deeply impassioned and imbued with strong (negative) emotions (Bonansinga 2020a, Salmela & von Scheve 2018). Although the nature of the relationship between emotionality and populist discourse has rarely been subject to systematic empirical scrutiny (Widmann 2021), this link has become one of the main focal points of research on political communication in recent years. Comprehensive research on the matter is thus crucial for gaining insight into the extent to which populist narratives and performances are rooted in and shaped by emotion, and by extension, how emotion may be utilized as a tool for building and maintaining political support (Nguyen 2019).
In the past, studies of populist communication have generally focused on its verbal elements, by analyzing the affective content of the words used in oral or written addresses of political leaders (Bobba 2019, Breeze 2019, 2020, Cochrane et al. 2022, Jacobs et al. 2020, Martella & Bracciale 2022, Widmann 2021). While advantageous in its own way, this type of analysis can pose a challenge to researchers primarily due to language barriers and the fact that written text usually contains only a small amount of words that have an inherent affective connotation (Widmann 2021). At the same time, currently available sentiment analysis tools are predominantly based on and developed for the English language, while the output for other languages is often of lower quality (Mohammad 2016). These circumstances therefore limit the possibility of result generalization and cross-cultural comparison.
Visual analyses, although insufficiently represented in the given field of research, may provide further insight into the characteristics of affective populist communication as they are not hindered by the same limitations as text-based studies (Moffitt 2022b) and allow for analysis of the visual rhetoric constructed and utilized by political leaders (Mendonça & Caetano 2021). Within the scope of visual research on political emotion display, two main schemes of analysis have shown to be especially valuable: the ethological approach (Masters et al. 1986) and the Facial Action Coding System (FACS) (Ekman & Friesen 1976).
Studies within the former largely focus on the analysis of political candidates’ non-verbal representations in news coverage, their displays during political campaigns and debates, as well as the impact of different forms of representation on viewers and success in political elections (Bucy et al. 2020, Bucy & Grabe 2007, 2008, Grabe & Bucy 2009, Sullivan & Masters 1988). The Facial Action Coding System, on the other hand, allows for a more in-depth analysis of the (micro)expressive behavior of political leaders during speeches and public performances. Due to its complexity, Ekman’s classification system has been applied in a limited number of studies (Stewart et al. 2009, 2015, Stewart & Ford Dowe 2013) but as the process of visual analysis gradually becomes automated (Joo et al. 2019) this research method is becoming more accessible to researchers from various fields of interest (Ash & Caliskan 2022, Bossetta & Schmøkel 2023).
Our study expands on the existing framework of FACS research on political emotion display and applies a computer vision approach for detecting emotions from the facial expressions of political leaders. Because it is relatively new, this approach faces various methodological challenges which we aim to address by following the recommendations outlined by Barrett et al. (2019), such as using big data techniques and automated detection in naturalistic settings, integrating traditional methods with machine learning, and sampling visual data across diverse contexts and cultures.
Since the introduction of the FER dataset (Goodfellow et al. 2013), different machine learning algorithms have been developed for solving the challenge of facial expression recognition, which essentially requires performing two tasks: (1) detection of facial expression, and (2) emotion recognition from the detected facial expressions. We develop a reproducible and principled computational workflow in Python for the analysis of video content of political leaders’ public performances. Our work relies on the existing Python library fer (Shenk et al. 2021) which bundles a pre-trained convolutional neural network (CNN) Keras models for face detection (Zhang et al. 2016a, based on) and emotion recognition (Arriaga et al. 2017, 2020).
We find that populist leaders on average express more negative emotions and that the emotionally neutral facial expression is less frequently detected in their public appearances compared to non-populist leaders. Overall, our work points towards a new direction in computational studies of the affective aspects of populist political communication and provides a roadmap for overcoming the limitations of the current approach in future studies.
2 Populism as performance
Defining populism represents a challenging task for researchers as there is barely any consensus on what such a definition should (or should not) include, and the concept itself becomes increasingly muddled as the term is used to describe a growing number of phenomena (Gidron & Bonikowski 2013). Most conceptualizations of populism can, however, be classified (Rovira Kaltwasser et al. 2017) into one of three main approaches: (1) the ideational approach (Mudde 2004), (2) the discursive approach (Norris & Inglehart 2019, Ostiguy 2017) and (3) the political-strategic approach (Weyland 2017). For the purposes of this study, we adopt the discursive approach, which conceives populism primarily as a rhetorical style based on first-order principles pertaining to who should rule—the people, as “the only legitimate source of political and moral authority in a democracy”—but which says very little about the how (Norris & Inglehart 2019). Whereas the ideational approach claims populism is a “thin-centered” ideology (Mudde 2004) and emphasizes the interpretation of politics as “a Manichean struggle between a reified will of the people and a conspiring elite” (Hawkins & Kaltwasser 2018) (Hawkins and Rovira Kaltwasser, 2018), the discursive approach highlights the ways in which this interpretation is framed and communicated as an anti-establishment discourse (Aslanidis 2017). We find this to be the most comprehensible understanding of populism as it considers the operationalized elements of the populist communication logic (Engesser et al. 2017) and examines the rhetorical components implicit in the ideological definition (Aslanidis 2016).
From this perspective, the populist rhetoric has widely been characterized by the use of “dramatization, polarization, moralization, directness, ordinariness, colloquial and vulgar language” (Engesser et al. 2017), and low-order appeals within which leaders “use a language that includes slang or folksy expressions and metaphors, are more demonstrative in their bodily or facial expressions as well as in their demeanor, and display more raw, culturally popular tastes” (Ostiguy 2017). Because it represents a form of symbolic resistance to the more formalized and composed self-representation of mainstream pluralist leaders, i.e., the “establishment” and “political elite”, the populist performance  [^1] This does not imply that, on average, there is a tendency for Japanese leaders to be more neutral. In fact, in Japan, we observe the same tendency as seen in the entire sample — the speeches of Shinzo Abe (LDP) from the populist group have the highest average score of negative emotions (0.58), while Natsuo Yamaguchi (NKP) from the moderately pluralist group has the lowest average score (0.30). Conversely, the average scores of neutrality show the opposite trend. These results are not extreme compared to the sample distributions shown in 2.is subversive to the normative communication structures of political dialogue. As such, it appeals to “common sense” (Bucy et al. 2020) and claims to give a voice to “the ordinary people” (Inglehart & Norris 2016).
This mode of communication has shown to be highly effective as a means of garnering media attention and cultivating a dedicated base of supporters—especially in the new media environment—and populist leaders seem to be overall more adept at capitalizing on it than other, non-populist political figures (Bobba 2019, Davidson & Enos 2022). Since digital media hinges on the attention economy (Wells et al. 2020) and thus tends to favor “dark participation” (Larsson 2022), populist leaders and political parties may fare well online due to the transgressive nature of their self-representation and the simplification, emotionalization, and negativity inherent in populist messaging (Bucy et al. 2020, Engesser et al. 2017). Social media is a particularly convenient tool for the proliferation of populist messages and imagery as it allows political actors to circumvent the mainstream media, reach large audiences, and communicate and interact directly with ‘the people’ (Lalancette & Raynauld 2019, Moffitt 2022a). At the same time, online platforms allow for an instant audience response, as the “feedback loop between producer and consumer is much more intense and rapid than with television and radio (Finlayson 2022). This form of ‘mediatized populism’ (Mazzoleni 2014) allows for aspecific type of political performance, one which is largely unmitigated and relies heavily on portraying political leaders favorably to (online) audiences by framing them as expressive, dominant, and disruptive figures who are in tune with “the will of the people”.
In recent years, much attention has been paid to the strategies of (affective) populist communication online, especially through analyses of verbal content shared by populist leaders and parties on Twitter and Facebook (e.g. Bracciale & Martella 2017, Gründl 2022, Jacobs et al. 2020, Martella & Bracciale 2022, Widmann 2021). Given the rise of visual communication in politics and the rapid developments of Web 2.0 practices (Grabe & Bucy 2009, Larsson 2022, Messaris 2019), our primary aim was to expand the scope of research on the affective self-representation of populist leaders online by examining the characteristics of their non-verbal emotional communication. The facial expressions of political leaders are a particularly valuable area of research in this context, since digital media provide a breadth of material for analysis—well-lit, high-quality close-up shots of political leaders are now more prevalent than ever. We used this abundance of available data sources to our advantage and processed 220 YouTube videos uploaded to official accounts of political parties and their leaders, amounting to a total of 77 hours of video material. Our analysis examines potential differences in facial expressions between populist and non-populist political leaders during public performances, and more specifically, the characteristics of their negative emotional expressions and neutrality in emotional stance (i.e., absence of emotional expression).
3 Data & Method
To create our sample, we first had to differentiate and categorize political parties (and, by extension, their leading figures) as either populist or pluralist. For this purpose, we used data from the Global Party Survey (GPS) (Norris 2020), an expert survey from 2019 whichoffers an overview of key ideological and issue positions for 1043 parties from 163 countries worldwide. While expert surveys certainly have their limitations (Wiesehomeier 2018) and may not be able to fully capture the nuances of populist rhetoric, this database remains the most comprehensive comparative overview of political parties globally. Among other data, such as ideological framework and party size, the survey also offers an estimate of the degree to which existing parties employ populist rhetoric. In the questionnaire design, populist rhetoric is defined as “language which typically challenges the legitimacy of established political institutions and emphasizes that the will of the people should prevail” (Norris 2020), whereas pluralist rhetoric “rejects these ideas, believing that elected leaders should govern, constrained by minority rights, bargaining and compromise, as well as checks and balances on executive power” (Norris 2020). The populist rhetoric scale in (V8 in GPS dataset) is based on expert assessments of rhetorical style given these parameters. The variable Type_Populism categorizes this scale into four ordinal groups: strongly pluralist, moderately pluralist, moderately populist, and strongly populist. We also use the “Type_Values” variable in supplementary analysis to examine ideological differences in expression of emotions (see Supplement section E – Ideological differences).
From the GPS dataset, fifteen countries with competitive popular elections were chosen: Argentina, Australia, Brazil, Croatia, France, Hungary, India, Italy, Japan, the Netherlands, Serbia, South Africa, Turkey, the United Kingdom, and the United States. This sample, although limited, provides somewhat broad coverage of various cultures and political systems globally. Three parties with varying degrees of populist rhetoric—strongly populist, strongly pluralist, and either moderately populist or moderately pluralist—were selected from each country. We primarily focused on minor and major political parties as defined by the GPS variable categorizing parties based on their share of the vote in contests for the lower house of the national parliament. Fringe parties (less than 3% of the vote) were generally excluded from the sample since their audience reach and influence is considerably less than that of their larger counterparts competing at the same scale in their respective countries. Likewise, due to this criteria, only two parties from the US were included in the sample. Table LABEL:tab:si-lead from the Supplementary Information contains information about the selected parties and their leaders.
We then identified one leader per party who has held a prominent political position over the past decade (up to 2019, covered by the GPS), and found their official YouTube channel or their party’s official YouTube channel. From these channels, five videos were selected for each leader based on the following criteria:
  1. (1)
    View count. Videos that reached a larger audience were given priority over videos that did not perform comparatively as well on the platform. View count was used as a filtering mechanism to single out videos with large and potentially diverse audiences, because the most viewed videos on the official YouTube channels of major political parties usually refer to topics of wider public interest. This is partially due to the “popularity bias” on YouTube, meaning it has a tendency to recommend political content that aligns with the interests of the majority (Heuer et al. 2021).
  2. (2)
    Unmediated representation. Only direct and unmediated representations of the given political leaders were included, i.e., videos which were not in any way previously modified by third parties such as news outlets. These consisted mainly of speeches, press conferences, ads, and promotional videos. This was done to ensure that the given results would reflect the leaders’ self-representation, rather than their representation in news coverage.
  3. (3)
    Consistent framing. A preference was given to videos where the face of the leader was in focus for an extended period of time and which did not frequently switch between disparate camera angles. This allowed for an overall consistent frame of reference for each given video.
  4. (4)
    Video quality. High-quality, well-lit videos were selected whenever possible so as to improve analysis accuracy.
This selection process resulted in a sample of 220 videos, representing 44 political leaders from 15 different countries. The sample is not gender-balanced as only five female leaders are present, which reflects the global underrepresentation of women in leadership positions within major political parties. As a result, our analysis does not account for gender, which we acknowledge as a limitation of our approach. Additionally, because the sample is limited by the timeframe of the currently available GPS dataset, it hasn’t captured more recent global developments on the populist-pluralist spectrum (e.g. Giorgia Meloni and the Brothers of Italy movement or the Servant of the People party in Ukraine are not present in the dataset). Given these limitations, our sample is conceived as a diverse snapshot of the state of populist and pluralist rhetoric in the second decade of the 21st century and does not capture the full scope or evolution of populism today.
We processed each video in the sample by extracting 300 frames uniformly distributed over the length of the video, which were the basis for our main analysis. To improve the robustness of the overall approach, we also included an alternative frame processing approach in the supplementary material, based on processing every 50th frame of the video. Given that the majority of the YouTube videos have 24–30 frames per second, this is equivalent to selecting a frame every 1.67–2.00 seconds, which is a reasonable time window if we expect to capture potential changes in facial expression. The average number of frames per video extracted using this approach was 273.03, which makes the resulting time-series of emotion scores comparable in length with the first approach.
For each selected frame of every video in the sample the following procedure is applied. First, face detection is applied to the extract frames using a multi-task cascaded framework based on convolutional neural networks (CNN) (Zhang et al. 2016b). This framework is suitable for detecting faces in videos because it performs simultaneous face detection and alignment, which is important as we cannot expect that political leaders will be looking straight at the camera in each extracted video frame. After face detection was applied to all frames, there were videos in which multiple faces were detected throughout the frame. These videos were manually checked to confirm which of the detected faces corresponds to the selected political leader.
The second step is emotion detection within the verified detected face box. We relied on implementing the mini-Xception architecture of CNN (Arriaga et al. 2017) aimed at detecting 6 basic facial expressions of emotion, as well as a neutral expression. Similar to the face detection approach outlined above, this implementation is aimed at real-time emotion detection from video input for robotic applications. This model achieves 66% accuracy (see Arriaga et al. 2017, p. 3 for more details) on the FER2013 dataset (Goodfellow et al. 2013) which is a standard benchmark for emotion detection. The dataset consists of 30,000 small images (48x48 pixels) showing diverse facial expressions of different emotions.  [^2] FER2013 was constructed using Google Images API queries for emotion-related keywords like “blissful” and “enraged”, combined with queries introducing diversity in gender, age and ethnicity represented in the images. Images were then labeled by human coders (Goodfellow et al. 2013, pp. 3-4). Unfortunately, no information is provided by the authors of the dataset regarding the number of human coders nor their background, or the labeling methodology applied to extract training labels for every image included in the FER2013 dataset. Benchmarks based on the FER2013 dataset should be taken with caution, especially in social science applications, because the dataset construction and validation documentation lacks crucial information about inter-coder reliability, coding procedures, and potential biases in the image selection process. Each emotional expression has approximately 6000 images, except disgust, which is only present in 600 images. According to Goodfellow et al. (2013), human classification accuracy is in the range of $65\%\pm 5\%$, but no additional information about the sample size, number of coders, or detection procedure is given.
In order to provide additional information and to validate the output of the machine learning approach, we conducted a small image annotation study where 5 coders annotated a sample of 514 randomly selected frames extracted from 220 videos. Coders received basic training in emotion recognition and, due to limited resources, answered only two questions about every image: (1) Can you clearly see the face of a politician in this image? (Yes/No), (2) Is this face expressing positive emotion, negative emotion, or is it a neutral expression? (“Positive emotion”, “Negative emotion”, “Neutral expression”). The majority vote label was set as the “human-coded label” and compared with the output of the machine learning algorithm, whose label is defined based on the highest score out of the sum of positive emotions, negative emotions, or the neutrality score. A detailed description and the results of the data annotation study can be found in Supplement F.
The result of the emotion detection procedure on a single video frame is an array of 7 emotion scores, each representing the predicted proportion of an expressed emotion. Each score is matched with one of 7 labels: “angry”, “disgust”, “fear”, “happy”, “sad”, “surprise” and “neutral”. The lowest possible score indicating the absence of a particular emotion is 0, while the highest score, which indicates a “clean” expression of a single emotion, is 1. In a statistical sense, the resulting data are compositional as they do not carry absolute, but rather relative information about the expressed emotions. This data is organized as a data frame consisting of 7 columns and approximately 300 rows, storing emotional scores for each selected frame of the video.
Because we were primarily interested in neutral emotional expressions and expressions of negative emotions, we created an additional variable which sums the scores of negative emotions: anger, fear, disgust, and sadness. Figure 1 shows examples of faces with a high detected score of negative emotion and a high score of neutral expression.
Figure 1: Example frames showing a high detected score of negative emotions (top: angry 0.78, disgust 0.00, fear 0.09, happy 0.00, sad 0.07, surprise 0.00, neutral 0.04) and a high detected score of neutral expression (bottom: angry 0.07, disgust 0.00, fear 0.04, happy 0.00, sad 0.02, surprise 0.08, neutral 0.79).
The entire analysis was performed in Python and the code is publicly available in the GitHub repository: https://github.com/atomashevic/face-of-populism. The code is organized into three sections. The data processing part of the code can be used to reproduce emotion score data files from the list of YouTube URLs for the videos included in the sample. It can easily be modified to expand the sample to a larger number of videos. Face and emotion detection relies heavily on the fer Python library and that part of the code is presented in the Jupyter notebook. Finally, we have data analysis scripts for performing data wrangling and statistical tests as well as producing Figures 2, 3, and 4.
4 Results
In the entire sample, there are 203 successfully processed videos using the approach of selecting 300 frames uniformly distributed across the entire length of the video clip. For 17 videos, multiple faces were detected without consistent classification of each person as a unique face, and these videos were discarded from the analysis. Pluralist leaders are present in 96 videos (31 strongly pluralist and 65 moderately pluralist) and populist leaders are present in the remaining 107 videos (69 strongly populist and 38 moderately populist). For every video, the emotion recognition procedure returns the score of the 6 detected emotions (anger, disgust, fear, happiness, sadness, surprise) as well as the score of the neutral expression for each of its frames. These scores sum up to 1 and they can be viewed as proportions of each emotion present in the compound facial expression detected in the frame.
Human image annotation was performed on a sample of 514 images extracted from 220 videos. Agreement between human coders ranged from 30.3% to 69.8%, while Cohen's Kappa coefficient varied from 0.15 to 0.462. Overall, Fleiss' Kappa is 0.231, which suggests fair agreement between coders but is also indicative of how difficult the task of recognizing the emotional expression in a single, randomly selected video frame is. When only selecting cases where a majority vote label exists (458 images), agreement between that label and the result of the machine learning classification is 55.2%, with a Cohen’s Kappa of 0.311 and F1 score of 0.502. This suggests moderate agreement and performance of the machine learning model relative to human consensus. When only selecting cases where the majority of coders and the model agree that a face is clearly visible (present) in the image (356 cases), agreement is 59.8% (Kappa 0.320, F1 score 0.530).
For each video, we calculated the mean score of each emotional state for all frames, as well as the mean of neutral expression, and the mean of negative emotional expression for the entire video. Descriptive statistics of these values are presented in Table 1.
Mean SD Min. Median Max.
Anger 0.227 0.134 0.034 0.200 0.726
Disgust 0.003 0.008 0.000 0.001 0.066
Fear 0.127 0.091 0.014 0.099 0.548
Happiness 0.077 0.083 0.000 0.051 0.414
Sadness 0.206 0.110 0.023 0.181 0.568
Surprise 0.031 0.037 0.000 0.020 0.273
Neutral 0.328 0.169 0.020 0.300 0.807
Negative emotions 0.563 0.185 0.156 0.570 0.964
Table 1: Descriptive statistics for extracted emotions for the entire sample, $N=203$ videos
Results suggest that negative emotions are more frequently expressed in the entire sample, with anger (0.227) as the emotion having the highest average score, followed by sadness (0.206) and fear (0.127). As a consequence, the average proportion of negative emotions is greater than 0.5, which means that on average negative emotions are detected more frequently than positive emotions (e.g., happiness) and neutral expressions combined. There are clear issues with the detection of disgust and surprise, where a low maximum level suggests low positive predictive value for these emotions. However, our focus is on differences in the expression of negative emotions and in scores of neutral states. Aggregating negative emotions enables us to surpass the false positive issues within the group of negative emotions (e.g., expression of fear being classified as sadness, disgust as anger, Khaireddin & Chen 2021) and minimizes the impact of the low positive predictive value of disgust and surprise.
Looking at the maximum values of average detected scores, the largest average scores of specific negative emotions were found in videos of populist leaders. The highest proportion of anger was detected in a video of Andrej Plenković (HDZ, Croatia), disgust in a video of Narendra Modi (BJP, India), fear in a video of Saša Radulović (DJB, Serbia) and sadness in a video of Donald Trump (GOP, USA). The video with the highest aggregate proportion of negative emotions, where one of the negative emotions is dominant in almost every analyzed frame, is the video titled “Special Message from President Trump”, published on December 23, 2020. On the other hand, the highest average score of neutral expression is found in a video of Yukio Edano (CDP, Japan), whose party belongs, in contrast with previous examples, to the moderately pluralist group. [^3] This does not imply that, on average, there is a tendency for Japanese leaders to be more neutral. In fact, in Japan, we observe the same tendency as seen in the entire sample—the speeches of Shinzo Abe (LDP) from the populist group have the highest average score of negative emotions (0.58), while Natsuo Yamaguchi (NKP) from the moderately pluralist group has the lowest average score (0.30). Conversely, the average scores of neutrality show the opposite trend. These results are not extreme compared to the sample distributions shown in Figure 2.
To further examine differences between pluralist and populist leaders, we used the GPS measure of populist rhetoric in a binary way to split the sample into groups of pluralist (Party Populism 1-2) and populist (Party Populism 3-4) leaders. The distribution of the average scores of negative emotion and neutral expression is shown in Figure 2.
Figure 2: Distribution of average scores of negative emotions (left) and neutral expression (right) of videos featuring populist and pluralist leaders.
Overall, all average values have a large spread in both groups, with pluralist leaders having several videos with less than a 0.2 average negative emotion score and more than a 0.75 average score of neutral expression. In terms of central tendency, we see denser grouping of pluralist leaders between the 0.4 and 0.6 negative emotions score, while populists tend to cluster in areas between 0.6 and 0.8. In terms of neutral expression, we have a thicker cluster of points for populists in the range between 0.2 and 0.4. Figure 1 suggests that, for populist leaders, we are more likely to observe lower average scores of neutral expression and higher average scores of negative emotion. The reverse tendency can be observed in the case of pluralist leaders. Negative emotions are dominant in 73.8% of populist videos, compared to 45.8% of pluralist videos.
The mean average score of negative emotions in the populist group is 0.616 (95% CI 0.584–0.649), compared to the 0.500 mean average score for pluralists (95% CI 0.464–0.537). The difference between the two groups is statistically significant ($t=4.691$, $p=0.000005$, $d=0.66$ moderate effect size). This finding suggests that pluralists generally tend to have an average score of negative emotions around 0.5, which roughly translates to the situation where negative emotion is the prevailing expression in one half of the video (in fact, that percentage is 46.1% compared to 66.7% for populists; see Table 3 in the Supplement). On the other hand, for populists, more than one half of each video where they appear is dominated by negative emotions.
Figure 3: Raincloud plot showing the distribution of average scores of negative emotions across four degrees of populism. Every data point is presented as a colored dot and above them boxplots and density plots are displayed. The dashed vertical bar indicates the score of 0.5.
The mean average score of neutral expression for populists is 0.289 (95% CI 0.260–0.318), compared to the 0.374 mean for pluralists (95% CI 0.338–0.409). While this difference in means is statistically significant ($t=-3.636$, $p=0.00035$, $d=-0.515$ moderate effect size), it is smaller than the one observed in the case of negative emotions.
To investigate the robustness of these differences, we took into account the differences between moderate and strong pluralists/populists provided by the GPS. The distributions of average scores in these four groups are compared and shown in Figures 3 and 4.
Figure 4: Raincloud plot showing the distribution of average scores of neutral expression across four degrees of populism. Every data point is presented as a colored dot and above them boxplots and density plots are displayed. The dashed vertical bar indicates the score of 0.5.
In the case of negative emotions (Figure 3), we observe four flat distributions with thick tails. However, the median point of all distributions progressively moves towards more negative emotions as we move from strong pluralists to strong populists. What is apparent in these differences is the low frequency of videos featuring populist leaders where the average score of negative emotions is below 0.5. It is worth pointing out that examples of extremely high scores of negative emotions are present also in pluralist groups.
Similarly, in the case of neutral expression (Figure 4), we have flat distributions (except in the case of moderate populists) with long right tails containing videos of leaders who maintain a neutral expression throughout most of their recorded appearance. Again, median values move progressively towards lower neutrality as we move from pluralists towards populists, except that no large difference can be seen when comparing strong and moderate populists. In the case of strong populists, however, we have a larger number of videos where the neutral expression is almost completely absent.
Overall, the results of one-way ANOVA suggest that the differences between the group means are statistically significant both in the case of negative emotions ($F=7.109$ $p=0.000152$), with strong effect size measured by eta-squared (0.104), and neutral expression ($F=5.625$ $p=0.001038$), with moderate effect size (0.084). Results from Tukey HSD post-hoc tests reveal significant differences between leaders coming from parties employing strong populist rhetoric and leaders coming from pluralist parties (mean difference $0.1559$ $p=0.0005$ in comparison with strong pluralists, and mean difference $0.1148$ $p=0.0029$ in comparison with moderate pluralists). For neutral expression we find the opposite case, as strong pluralists differ from both groups of populists (mean difference $0.135$, $p=0.0056$ for moderate populists, and mean difference $0.129$ $p=0.0032$ for strong populists).
Overall, these results reveal differences in the average scores of negative emotion and neutrality between populist and pluralist leaders on a modest sample. These differences are more pronounced when we focus on candidates coming from parties who strongly rely on either pluralist or populist rhetoric, with strong pluralists on average favoring more neutral expressions, while strong populists more frequently express negative emotions—meaning that, for the majority of their public appearance, we can observe facial expressions of negative emotions.
5 Discussion
The present study examines facial emotional expressions of populist and non-populist political leaders across fifteen countries using a deep-learning-based computer-vision algorithm to investigate differences in their expressions of negative emotions (anger, disgust, fear, and sadness) and neutral stances, i.e., facial expressions which lack emotional disposition. Our findings indicate that populist leaders tend to express negative emotions more frequently than non-populist leaders, and conversely, that neutral expressions are less common in their public performances than they are in those of more conventional, pluralist political figures. The significance and the effect size of these differences are higher compared to differences between ideological groups (see Supplement E). These findings align with the theoretical framework on the characteristics of affective populist communication (Bonansinga 2020b, Salmela & von Scheve 2018, Engesser et al. 2017), as well as with previous, mostly text-based research which indicates that populist leaders employ language which is associated with negative emotion at a higher rate than mainstream, pluralist politicians (Breeze 2019, Caiani & Di Cocco 2023, Macagno 2022, Martella & Bracciale 2022, Widmann 2021).
Overall, the results provide new evidence of a style of affective, non-verbal communication more common to political leaders whose rhetoric is populist-leaning, and thus further substantiate the idea that populism is performative in nature Bucy et al. (2020), Moffitt (2022b). As such, because it is subversive in relation to established political conventions, the populist performance is especially well-suited for the environment of social media where the principle of “any press is good press” often applies—that is, the quantity of audience attention bears more weight than its quality (Wells et al. 2020). On one hand, the populist use of negative emotional displays in video may therefore be a mechanism through which they maintain relevance, considering that controversial content online ordinarily generates more engagement among consumers (Larsson 2022). On the other, populists’ facial expressions can provide affective heuristics for audiences, influencing their evaluation of candidates’ political competence (Boussalis & Coan 2021) and allowing them to quickly reduce complex political issues into simple, emotional interpretations by visually framing them as negative. Further analysis of the role of non-verbal cues in political communication—such as facial expressions of emotion—would be beneficial to gaining a better understanding of the populist performance as a whole.
Apart from these findings, the present study introduces an open-source workflow for the analysis of political videos and provides a new open dataset of time-series emotion scores, as well as tools for the annotation of images in R and a new dataset of human labelled images which can be used to benchmark different machine learning approaches to facial expression recognition. Overall, these contributions aim to make large-scale computational studies of facial expressions of emotion more accessible and rigorous in the future.
Given the exploratory nature of the study and its reliance on machine learning algorithms for emotional detection, there are several notable limitations with significant impact on the results presented in this paper. There is a group of limitations inherent to the way we constructed the sample, and another related to the machine learning approach to detecting emotions from facial expressions.
Firstly, the sample of YouTube videos is not globally representative and does not provide ample coverage of the global political spectrum in regard to the use of populist or pluralistic rhetoric. As we have shown in the supplementary analysis (see Supplement C: Country-level differences), the differences between populist and pluralist leaders are not consistent in all investigated countries. Larger research infrastructure is required to process a more substantial number of videos, leaders and countries and the current computational toolset needs to be extended to enable seamless automatic processing of videos. This also relates back to the gender imbalance of the current sample. While it is reasonable to hypothesize that gender influences emotional expression among political leaders (Boussalis et al. 2021), in thiscase, including more women without significantly increasing the overall size of the sample and the number of leaders would mean focusing only on countries where there is at least one female leader of a non-fringe political party. This creates the risk of introducing bias towards Western countries in the sample and disregarding the global scope.
Secondly, a key component of our analysis is the measure of populist rhetoric supplied by the Global Party Survey database (Norris 2020) which significantly limits the timeframe of the analysis (no leaders, parties and videos after 2019 were taken into consideration), but also imposes the inherent limits of an expert survey. While the GPS is the only available source of multi-dimensional measurements of populist rhetoric on a global scale, the experts surveyed may operate within different understandings of what populist rhetoric is. This is particularly challenging in developing countries, where fewer experts answered the survey, as well as in autocratic regimes where experts may not have been able to provide a fair assessment of the (ruling) political parties (or refused to do so). By focusing on large countries with free elections, we tried to mitigate some of the risks, but there is the possibility that some measurements used in this study reflect expert biases, especially since we used a single, aggregate measure of populist rhetoric. Future studies should aim for more granularity and robustness in the measurement of populist rhetoric, which may be obtained by following a text-as-data approach and obtaining the measures using computational and machine learning tools, where expert survey scores provide external validation (Jankowski & Huber 2023).
Furthermore, since we lacked the resources necessary to process every political speech from each party’s official YouTube channel, the first criterion of selection is the view count of the videos. Our expectation is that this metric serves as a proxy for the relevance of the speech content to a wider audience over a longer period of time. However, since we have no information about the audience, it is possible we are observing a demand-side effect, where audiences who consume populist videos might show a stronger preference for those that prominently feature negative emotions. This could introduce a specific bias in the sample.
One of the major limitations of our current workflow is that the video-processing algorithm does not focus on one singular face and does not consistently label the same faces between frame jumps, which requires researchers to select videos with consistent framing and a small number of different faces. Solving this problem will allow for faster and more reliable collection of larger samples.
The next major limitation is a result of the machine learning model deployed for emotion recognition. We conducted a small validation study to compare the performance of our model with the consensus of 5 human coders. Although results show evidence of moderate agreement between humans and the machine learning model, there is high inconsistency between human coders even for the simple task of assessing whether a given facial expression shows positive emotion, negative emotion, or a neutral expression. There is an urgent need for large-scale image annotation studies with expert coders if we are to assess the performance and accuracy of FER models in the future, especially on difficult “in the wild” image datasets such as ours. However, it is worth noting that in our main analysis we focus on averages across 300 frames of each video, making our analysis robust to misclassification of a smaller number of individual frames. That being said, we believe that fair agreement between the human majority vote and machine learning model does not compromise the validity of our findings.
Overall confidence in the presented results would surely be higher if we based our analysis on a computer-vision approach which was shown to have higher accuracy, at least on FER2013 data. One of the possible alternatives would be to follow recent work (Bossetta & Schmøkel 2023, Schmøkel & Bossetta 2022) and use proprietary API such as Amazon’s Rekognition. Nonetheless, the aim of this paper was to introduce an open-source approach focusing on free computing resources and the deployment of machine learning algorithms on systems that are accessible to individual researchers in social sciences, including graduate students without significant institutional support for computational work.
This study was focused on exploring basic mean-level differences between the given groups. It is worth noting that the full output of our emotion detection workflow includes time series data which are suitable for multivariate analysis. Even though time series are compositional in nature, they can be transformed (Greenacre 2021) so that traditional time series methods can be applied to explore affective dynamics (Ryan et al. 2023), find additional differences in relation to other factors (such as type of public address, the proximity of elections, ideology,etc.) and cluster political leaders according to patterns of correlations in time-series. Furthermore, multivariate change point detection methods (James & Matteson 2015) could be used to identify the key moments in speech where significant changes in emotional expression occur, which can further be linked to textual analyses of transcripts. This would allow researchers to identify the rhetorical elements which trigger switches in the emotional dynamics of leaders’ political addresses.
As computational studies of political communication rapidly embrace machine learning tools and algorithms, this study offers an extension of an open-source toolbox for the analysis of affective non-verbal political communication. Analyses presented in this paper are based on a novel dataset which is only a starting point for further studies, and which will hopefully expand in the future to encompass a more diverse sample of political leaders in videos obtained from different platforms. Gaining this kind of insight into the visual aspects of emotional political communication, coupled with the existing analyses of its verbal elements, holds the potential for a more comprehensive understanding of how leaders convey political messages, persuade audiences, and present themselves to the public using digital media platforms.
6 Data availability
Data files are stored as CSV files. Processed data used for both the main and supplementary analyses can be found at: https://github.com/atomashevic/face-of-populism. Repository contains a CSV file with a list of URLs of all YouTube videos used in the analysis as well as Python code that can be used to reproduce the CSV files with processed data. CSV file extracted from GPS data is available in the repistory and the original GPS dataset is available at: https://dataverse.harvard.edu/dataverse/GlobalPartySurvey.
Code used for image annotation and the resulting dataset can be found at: https://github.com/atomashevic/fop-annotation.
7 Acknowledgements
The authors did not receive any financial support for this research. The annotation of images was made possible thanks to the dedicated work and effort of our undergraduate students, A. Nadj, M. Milanko, and A. Mišić, who volunteered as coders.
Appendix A Supplementary: Alternative frame processing approach
In our main analysis we presented the results based on the frame-processing approach of selecting 300 uniformly distributed frames from each video. An alternative, which yields roughly the same number of frames per video (273.03 on average), is to select and process every 50th frame of each given video. With this approach, we arrive at 209 successfully processed videos—six more than in the main analysis, because we find more consistent face selection in several videos when the algorithm receives this type of input. For shorter videos, this reduces the probability that, when selecting exactly 300 frames, a significant proportion of these frames captures cutaways to other people or the audience instead of the leader delivering their speech. Pluralist leaders are present in 101 videos (31 strongly pluralist and 70 moderately pluralist) and populist leaders are present in the remaining 108 videos (70 strongly populist and 38 moderately populist). Descriptive statistics for emotion variables obtained from this frame processing approach are presented in Table 2.
Mean SD Min. Median Max.
Anger 0.223 0.130 0.031 0.195 0.728
Disgust 0.003 0.006 0.000 0.001 0.051
Fear 0.122 0.087 0.015 0.093 0.530
Happiness 0.084 0.099 0.001 0.053 0.594
Sadness 0.204 0.109 0.021 0.183 0.573
Surprise 0.030 0.037 0.000 0.018 0.275
Neutral 0.332 0.167 0.021 0.309 0.812
Negative emotions 0.552 0.182 0.147 0.558 0.965
Table 2: Descriptive statistics for extracted emotions for the entire sample, $N=209$ videos
These results are quite similar to those presented in Table 1, with no major deviations in terms of central tendency or variability of the 8 variables. Anger, sadness, and fear are the most common emotions and negative emotions have an average score higher than 0.5.
Again, we observe statistically significant differences in the means of negative emotions and neutral expression between populists and pluralists (negative emotions: $t=3.742$, $p=0.00024$; neutral expression: $t=-3.222,p=0.00148$). The sample mean differences are smaller than in the main analysis, but with no changes in the statistical significance (negative: 0.116 in main analysis, 0.091 alternative; neutral: -0.075 in main analysis, - 0.074 in alternative). The distribution of average scores is shown in Figure 5.
Figure 5: Distribution of average scores of negative emotions (left) and neutral expression (right) of videos featuring populist and pluralist leaders.
Figures 6 and 7 illustrate the differences in the distribution of mean emotion scores for four subsamples according to the degree of pluralist/populist rhetoric. The results of one-way ANOVA remain in accordance with the main analysis, but with weaker effect size in the case of negative emotions. The differences between the group means are statistically significant both in the case of negative emotions ($F=4.886$, $p=0.002712$) with moderate effect size measured by eta-squared (0.073), and neutral expression ($F=5.180$ $p=0.001846$) with moderate effect size (0.077).
Figure 6: Raincloud plot showing the distribution of average scores of negative emotions across four degrees of populism. Every data point is presented as a colored dot and above them boxplots and density plots are displayed. The dashed vertical bar indicates the score of 0.5
Figure 7: Raincloud plot showing the distribution of average scores of neutral expression across four degrees of populism. Every data point is presented as a colored dot and above them boxplots and density plots are displayed. The dashed vertical bar indicates the score of 0.5.
The results from Tukey HSD post-hoc tests reveal significant differences between leaders coming from parties employing strong populist rhetoric and leaders coming from pluralist parties (mean difference 0.1298 $p=0.0057$ in comparison with strong pluralists, and mean difference 0.0961 $p=0.0166$ in comparison with moderate pluralists). For neutral expression we again find the opposite case, as strong pluralists differ from both groups of populists (mean difference 0.138 $p=0.0037$ for moderate populists, and mean difference 0.1131 $p=0.0122$ for strong populists).
Overall, we conclude that using an alternative frame processing approach does not change the outcomes of statistical tests, but the reduced effect size in the case of negative emotions and overall shrinkage of the absolute differences between groups should certainly be noted.
Appendix B Supplementary: Dominant emotions per frame
Another way of conceptualizing differences in emotional expression is to look at the proportions of frames in which a certain emotion or type of emotional expression is dominant among the groups (Table 3).
[…]
Main analysis
Pluralists 0.162 0.001 0.094 0.072 0.123 0.019 0.428 0.461
Populists 0.262 0.004 0.084 0.062 0.234 0.010 0.345 0.667
Supplementary
Pluralists 0.176 0.000 0.096 0.073 0.132 0.020 0.438 0.484
Populists 0.248 0.000 0.080 0.076 0.225 0.011 0.358 0.640
Table 3: Mean proportion of all frames in which emotion or group of emotions has the highest score, and mean proportion of frames in which negative emotions have a sum score $\geq 0.5$, per category of leader and split by frame processing approach.
Appendix C Supplement C: Country-level comparison
The sample of YouTube videos and leaders who appear in them is not representative for any country included. We have selected only a few political parties and deliberately excluded small parties from the analysis. Nonetheless, here we provide a brief overview of raw and mean scores of negative emotions and neutral expression for every country in the sample, split by populist and pluralist leaders. Both Figure 8 and 9 show that the differences between populists and pluralists are not always consistent between the countries.
Figure 8: Mean and raw scores of negative emotions for every country (ISO country codes on x-axis) in the sample and for populist and pluralist leaders. Transparent points are raw scores, thicker points are means for every group.
For example, in the case of Serbia, Croatia and South Africa, pluralists show more negative emotions (average score $>0.5$ in all three cases). On the other hand, in Serbia, Croatia, Australia, the Netherlands and Brazil, we have cases where populists are on average more neutral in their public appearances.
Figure 9: Mean and raw scores of neutral expression for every country (ISO country codes on x-axis) in the sample and for populist and pluralist leaders. Transparent points are raw scores, thicker points are means for every group.
Larger samples with more coverage of political parties and movements are needed in order to investigate whether these differences and inconsistencies are consequences of sampling variability or true effects.
Appendix D Supplementary: Political leaders
Table 4: Information about political leaders present in the YouTube video dataset, the parties they are affiliated with, party or coalition acronyms from the GPS dataset, and populism and ideology categories from the GPS dataset (TYPE_POPULISM and TYPE_VALUES variables)
Country Leader Party Acronym Populism Category Ideology Category
Argentina Alberto Fernández Justicialist Party PJ Strongly populist Right Conservative
Mauricio Macri Republican Proposal PRO-UCR (coalition) Moderately pluralist Right Conservative
Sergio Massa Renewal Front 1País (coalition) Moderately populist Right Conservative
Australia Warren Truss National Party of Australia NAT Moderately populist Right Conservative
Bill Shorten Australian Labor Party ALP Strongly pluralist Left Liberal
Scott Morrison Liberal Party of Australia LP Moderately pluralist Right Conservative
Brazil Simone Tebet Brazilian Democratic Movement MDB Moderately pluralist Right Conservative
Luiz Inácio Lula da Silva Workers’ Party PT Moderately pluralist Left Liberal
Jair Bolsonaro Social Liberal Party PSL Strongly populist Right Conservative
Croatia Andrej Plenković Croatian Democratic Union HDZ Moderately populist Left Conservative
Zoran Milanović Social Democratic Party SPH (coalition) Moderately pluralist Left Liberal
Ivan Vilibor Sinčić Key of Croatia (formerly named Human Shield) OOP (coalition) Strongly populist Right Conservative
France Marine Le Pen National Rally FN Strongly populist Right Conservative
Emmanuel Macron Renaissance (formerly named En Marche) REM Moderately pluralist Right Liberal
François Bayrou Democratic Movement MoDem Strongly pluralist Right Liberal
Hungary Viktor Orbán Fidesz Fidesz Strongly populist Right Conservative
Ferenc Gyurcsány Democratic Coalition DK Strongly pluralist Right Liberal
Péter Jakab Jobbik Jobbik Strongly populist Right Conservative
India Narendra Modi Bharatiya Janata Party BJP Strongly populist Right Conservative
Sonia Gandhi Indian National Congress INC Moderately pluralist Left Liberal
Mamata Banerjee All India Trinamool Congress AITC Moderately populist Left Conservative
Italy Matteo Salvini Lega Nord CD (coalition) Strongly populist Right Conservative
Enrico Letta Democratic Party CCS (coalition) Strongly pluralist Left Liberal
Beppe Grillo Five Star Movement M5S Strongly populist Left Conservative
Japan Shinzo Abe Liberal Democratic Party LDP Moderately populist Right Conservative
Natsuo Yamaguchi Komeito NKP Strongly pluralist Left Conservative
Yukio Edano Constitutional Democratic Party of Japan CDP Moderately pluralist Left Liberal
Netherlands Geert Wilders Party for Freedom PVV Strongly populist Left Conservative
Alexander Pechtold Democrats 66 D66 Strongly pluralist Right Liberal
Mark Rutte People’s Party for Freedom and Democracy VVD Moderately pluralist Right Liberal
Serbia Aleksandar Vučić Serbian Progressive Party SNS Strongly populist Right Conservative
Zoran Lutovac Democratic Party DS Moderately pluralist Left Liberal
Saša Radulović Enough is Enough DJB Moderately pluralist Right Liberal
South Africa Julius Malema Economic Freedom Fighters EFF Strongly populist Left Conservative
Jacob Zuma African National Congress ANC Moderately populist Left Conservative
John Steenhuisen Democratic Alliance DA Moderately pluralist Right Liberal
Turkey Recep Erdoğan The Justice and Development Party AKP Strongly populist Right Conservative
Selahattin Demirtaş Peoples’ Democratic Party HDP Strongly pluralist Left Liberal
Kemal Kılıçdaroğlu Republican People’s Party SHP/CHP Moderately pluralist Left Liberal
UK Nigel Farage UK Independence Party UKIP Strongly populist Right Conservative
Jeremy Corbyn Labour Party Lab Moderately populist Left Conservative
Boris Johnson Conservative Party CON Moderately populist Right Conservative
USA Donald Trump Republican Party GOP Strongly populist Right Conservative
Hilary Clinton Democratic Party Dem Moderately pluralist Left Liberal
Appendix E Supplementary: Ideological differences
We examined differences in the expression of negative emotions and neutrality of expression between different ideological groups according to the GPS variable Type_Values, which combines two categorizations of political parties: economic values (Left = pro-state or Right = pro-market) and social values (Liberal or Conservative). The results of independent samples t-test are shown in the Table 5.
Negative emotions Neutral expression
t p d t p d
Right - Left 1.447 0.151 0.203 -2.135 0.034 -0.304
Conservative - Liberal 1.715 0.088 0.214 -1.474 0.142 -0.207
Table 5: Results of the t-test for differences in mean negative emotions score and mean neutral expression between political leaders belonging to parties differing on Right VS Left and Conservative VS Liberal values according to the GPS dataset
Appendix F Supplementary: Image annotation
For the image annotation study, 514 frames were randomly selected from the sample of all frames extracted from the videos in the main study. Image annotation was performed in RStudio (Posit team 2024) using a Shiny (Chang et al. 2024) application built using a customized version of the taipan R package (Kobakian & O’Hara-Wild 2018). Figure 10 shows the main interface of the application used by coders. Responses from each of the 5 coders were exported from the application as CSV files.
Figure 10: Taipan interface showing the image from the sample (top left), short instructions (bottom left), and the questions about the image along with navigation buttons (top right).
After aggregating the results from all human coders, a majority vote principle was applied to obtain the final label. This means that at least three out of five coders had to apply the same label out of the four labels given (“No face”, “Positive”, “Negative”, “Neutral”) to the image in order for that label to be taken as the output of the human annotation process. In the case of the machine learning output, we took the highest (sum) score to obtain the final label: “Positive”, “Negative” or “Neutral”. If all scores had missing values in the case of the machine learning output, we assigned a “No face” label to that image.
Agreement between pairs of human coders and between any individual coder and machine learning output is not highly relevant to the aim of the study, however we show values of pairwise Cohen’s Kappa on Figure 11 to illustrate the difficulty of the task and the absence of high levels of agreement between human coders.
Figure 11: Pairwise Cohen’s Kappa values for agreement between pairs of human coders or pairs of human coders and the ML model.
[…]
Positive 0.5 0.25 0.23 0.02
Neutral 0.13 0.49 0.30 0.08
Negative 0.17 0.2 0.59 0.04
No Face 0.09 0.03 0.29 0.59
Precision 0.26 0.50 0.72 0.48
Recall 0.5 0.49 0.59 0.59
Table 6: Confusion matrix with row-wise ratios, precision and recall values for agreement between the majority vote of human coders and the ML model output.
In summary, when considering a 4-label problem where human labels are obtained by a majority vote, we have 458 cases with 55.2% agreement, Cohen’s Kappa of 0.31 and F1 score of 0.502. Table 6 provides additional information. The highest accuracy is recorded in the case of negative emotions and the lowest accuracy is in the case of positive emotions. This is likely due to the imbalance within emotion labels on which the ML model is trained.
Table 7 shows values for a subset of cases where the majority of humans and the ML model agree that there is a face clearly visible (present) in the image (3-label problem). Here, we have 365 cases with 59.83% agreement, Cohen’s Kappa of 0.32 and F1 score of 0.53.
[…]
Positive 0.51 0.26 0.23
Neutral 0.13 0.6 0.27
Negative 0.18 0.2 0.62
Precision 0.27 0.49 0.80
Recall 0.51 0.6 0.62
Table 7: Confusion matrix with row-wise ratios, precision and recall values for agreement between the majority vote of human coders and the ML model output in cases where there is agreement about a face being clearly visible (present)